OpenStudy (anonymous):

In triangle XYZ, the length of side XY is 25 mm and the length of side YZ is 36 mm. Which of the following could be the length of side XZ? A. 66 mm B. 9 mm C. 63 mm D. 41 mm

OpenStudy (defiance):

Hi.

OpenStudy (defiance):

Heres the equation you should use. \[a^{2} \times b ^{2} = c^{2}\]

OpenStudy (defiance):

Lets say XY is "a" (25) and YZ = b (36)

OpenStudy (defiance):

\[25^{2} \times 36^{2} = ?^{2}\]

OpenStudy (defiance):

I have a feel i messed up :D

OpenStudy (defiance):

@phi

OpenStudy (defiance):

@ganeshie8

OpenStudy (phi):

XY 25 mm and the length of side YZ is 36 if we imagine the two sides can fold, the smallest side would be

OpenStudy (phi):